### v1

@@ 6.4 Ablation Study
Effectiveness of multi-granular parsing data. We evaluate the text recognition and grounding performance on DocLocal4K [12]. As shown in Tab. 4, DOGE surpasses mPLUG-DocOwl-1.5 across all granularities. We also train a model called DOGE-, which is pre-trained without using our constructed multi-granular parsing data. Despite the domain gap between our parsing data and DocLocal4K due to their different construction methods, it is evident that incorporating our data improves text grounding and recognition accuracy across various text granularities. This validates the effectiveness of our multi-granular parsing data in enhancing basic text grounding and recognition capabilities.
Increasing the number of training image tiles leads to a significant improvement in high-resolution document understanding. As shown in rows 1 and 2 of Tab. 5, we evaluate model performance with varying numbers of training image tiles on two datasets: InfoVQA and the PDF Document subset of DOGE-Bench, both of which consist of high-resolution images. The results show that increasing the number of training image tiles from 9 to 16 significantly improves answer accuracy on both datasets. This validates the importance of providing sufficiently high-resolution input for effective document understanding in MLLMs.
Increasing the number of image tiles during inference is detrimental to grounding and referring. As shown in rows 2 and 3 of Tab. 5, we compare model performance with different numbers of image tiles during inference. The results show that increasing the number of tiles improves performance on general document QA tasks, with a 0.9% gain on InfoVQA. However, there is a notable decline in performance on the PDF Document subset of DOGE-Bench, including reduced accuracy for text answers and F1 scores for grounding performance. This suggests that the current bounding box modeling method is sensitive to out-of-domain input resolutions, highlighting the need for a more robust and generalizable bounding box modeling method for varying input resolutions.
@@ 7 Conclusion
In this paper, we introduce DOGE-Engine, a data construction pipeline for generating high-quality ground-and-refer data for fine-grained document understanding. Additionally, we construct DOGE-Bench as the first comprehensive benchmark for evaluating the document grounding and referring capabilities of MLLMs. Furthermore, leveraging data generated by our engine, we develop DOGE, a pioneering MLLM that integrates text grounding and referring abilities into the dialogue and reasoning process. Our results show that DOGE can achieve versatile document grounding and referring while achieving promising performance on traditional document understanding tasks. We hope our work will facilitate practical document AI assistants.
@@ B.2 Instruction Details
We add different response format prompts to different questions based on the format of the responses to help the model output corresponding results when users interact with it. For generated question and answer pairs, we combine different question and answer pairs with various response format prompts to obtain diverse grounded data. For Grounded Answering Data, we have 7 response format prompts to be added after the question, and the answer should be directly a simple text wrapped with “<ocr></ocr>” and followed by the coordinates wrapped with “<bbox></bbox>”. For Grounded Reasoning Data, we combine two random-chose prompts together to form the response format prompt, and the response should be such a sentence structure involves a segment of grounded reasoning followed by ”Answer: ” and a concise answer. For Grounded Open-ended Answering, we simply use the first part of the response format prompts for Grounded Reasoning, and the response should be a grounded reasoning sentence. For Plain text Answering, we add no response format prompt to keep consistent with the original question answering.
@@ Appendix C Qualitative Results
We present the inference results of DOGE for grounding, grounding-and-referring, and referring tasks. The specific coordinates of the annotations are omitted, and the grounded text is highlighted with colored boxes. The colors of the boxes within the document image correspond to the colors of the text boxes. For the grounding task, we present four examples. The first sample demonstrates that DOGE can perform fine-grained question answering on general document images with rich content and complex layouts, successfully grounding the corresponding information. The second sample illustrates the model’s excellent recognition and localization capabilities for diverse and small text in poster-type images. Additionally, we showcase two samples of grounding in chart-type images, which indicate that DOGE possesses a certain level of mathematical ability, enabling it to provide grounded reasoning during calculation, as well as the capability to estimate values in charts that lack textual annotations based on the axes. For the grounding-and-referring and referring tasks, DOGE is able to recognize the content of user-selected regions and provide reasonable grounded or plain textual reasoning and responses. DOGE exhibits robust fine-grained grounding and referring capabilities, allowing for reliable grounded reasoning and accommodating diverse user interactions, significantly enhancing the overall user experience.
@@ C.2 More Qualitative Results
C.2.1 DOGE-Bench Examples
As shown in Fig. 10, we present three inference examples of DOGE-Bench, demonstrating unlabeled value reading capability, grounding summarization capability, and referring summarization capability. The first figure is an example of reading unlabeled values. We add auxiliary mark boxes in red to the coordinate axes to facilitate reading the values, each box’s height represents 5. It can be observed that the model’s output values are accurate. The figure on the lower left is an example of grounding summarization for an entire document. DOGE is able to perceive the content of the entire document and perform grounding summarization effectively. The figure on the right is an example of referring summarization for a specific area. The model can also accomplish this task, thereby helping users improve reading efficiency.
Figure 10: DOGE’s inference results on DOGE-Bench.
@@ C.2.2 Other Examples
Strong generalization. As shown in the first row of examples, we feed DOGE with screenshots of the paper content and the specially shaped fan chart that the model does not process before. DOGE is able to correctly respond, demonstrating its strong generalization and usability in actual document reading scenarios.
Handwriting ground-and-refer ability. The middle sample showcases the model’s ability to recognize and grounding handwritten content. DOGE can fully understand this casual handwriting and provide grounded output. It is worth mentioning that our training data does not contain the grounding or referring tasks on such handwritten images.
Other capabilities. The bottom sample demonstrates an additional untrained capability of our model, specifically referring translation. DOGE is able to provide translations for the asked region. It is important to note that our training data do not include tasks similar to referring translation, even rarely includes languages other than English. Therefore, we believe that DOGE can effectively handle the relationship between regions and corresponding text, and seamlessly integrate the capabilities of large models with grounding and referring abilities.
@@ C.2.3 Failure Cases
Although DOGE demonstrates strong grounding and referring capabilities, there are still some shortcomings, as shown in Fig. 12.
Referring. The upper left figure illustrates an example of incorrect referring. It mistakenly associates the bounding box that should correspond to the text “INITIATIVE” in the question with the text “LOYALTY” below it. Additionally, its width is re-estimated according to the size corresponding to ”LOYALTY,” resulting in an incorrect answer.
Grounding. When encountering some unfamiliar text content such as tables in the upper right image, DOGE can understand the content, effectively identified the line breaks between “Doc” and “VQA” and merged them together, and provide correct answers, but the grounding boxes are inaccurate. There is also a issue of incomplete grounding content, such as the bottom sample, which often occurs when the content requiring grounding is interrupted or wrapped to the next line. Although this does not affect understanding, the text and bounding boxes provided by the model do not completely match.
@@
Figure 11: Other DOGE’s inference samples.
Figure 12: Failure cases of DOGE.
@@ D.1 Dataset Statistic
As shown in Tab. 6, we present the detailed composition of the DOGE dataset’s Multi-granular Parsing data and Instruction Tuning data. The Multi-granular Parsing data includes five granularities: word, phrase, line, paragraph, and full-page parsing. The Instruction-Tuning data comprises three types of grounded data: grounding referring, grounding-and-referring, and plain text Q&A. We provide the detailed data volume for each type.
@@
Table 6: The detailed composition of the DOGE dataset’s Multi-granular parsing data and Instruction-Tuning data.
@@ D.2 Construction Cost Details
In terms of multi-granular parsing data construction, DOGE-Engine can achieve boundary box annotation and content extraction on any document dataset with a similar rendering process. It can also scale up to a larger volume of PDF source files without additional manual costs. When constructing instruction fine-tuning data, our main cost lies in the API calls to GPT-4o. We use the version gpt-4o-2024-08-06. For poster and chart, our input is long text containing full page parsing data, while for PDF, the input is document images plus short prompt texts. During construction, we can generate multiple question and answer pairs for a single image simultaneously and use batch API requests to save costs. Ultimately, we can construct over 1,000 grounded question and answer pairs for an average cost of \$1, which is far more efficient and cost-effective than manual construction.
@@ Appendix E Training Data Details
For the pre-training data, we utilize DocStruct4M [12] along with our 2.1M multi-granular parsing data to enhance DOGE’s foundational grounding and referring capabilities.
@@
DOGE- 83.34 60.96 41.67 36.95 55.73 49.66 58.62 50.34 47.27 51.47 68.86
DOGE 92.94 94.05 86.52 86.08 89.9 73.57 86.48 76.85 71.8 77.17 95.24
Appendix F More Experiments
@@
Effectiveness of our pretraining data. To better assess base performance of pretrained model using our multi-granular parsing data, we further introduce our text&bbox localization test set DOGE-Local15k comprising four granularities (word, phrase, line, paragraph) across three categories of data: poster, chart, and general document, similar to DocLocal4k[12]. DOGE-, same as Tab. 4, is pre-trained without using our constructed multi-granular parsing data. For the chart data, we claim that the chart type annotations in DocStruct4M for during DOGE-’s pre-training are inconsistent with our grounding training objectives, which aims to match the text and the bbox. In the original data, the text corresponds to bounding boxes of bar/line charts. Therefore, these value is not referenceable, and we mark them in gray. Moreover, in the case where the granularity of data categories for charts is relatively singular, we directly reported the results for “ALL”.
@@
As shown in Tab. 8, after incorporating our multi-granular parsing data for pre-training, the model exhibits enhancements in text recognition and grounding tasks. Our improvements on DOGE-Local15k are significant, thereby laying a solid foundation for accurate document grounding interactions.

### v2

@@ 6.4 Ablation Study
Effectiveness of multi-granular parsing data. We evaluate the text recognition and localization performance on DocLocal4K [14]. We set our model pre-trained with only DocStruct4M[14] as our baseline. As shown in Tab. 4, with our multi-granular parsing data, the model achieves performance improvements at all granularities compared to the baseline. Despite the domain gap between our parsing data and DocLocal4K due to the different construction methods, it is evident that incorporating our data improves text grounding and recognition accuracy across various text granularities. This validates the effectiveness of our multi-granular parsing data in enhancing basic text localization and recognition capabilities.
Effectiveness of instruction-tuning data. Besides endowing the model with grounding and referring capabilities, instruction-tuning data further promotes the performance in document understanding. The model fine-tuned using data from DocOwl-1.5[14] serves as the baseline. As shown in Tab. 5, integrating our instruction-tuning data with DocOwl-1.5’s data leads to a 2.87% increase in ANLS on InfoVQA and a 21.35% increase in CIDEr on VisualMRC.
@@ 7 Conclusion
In this paper, we introduce DOGR-Engine, a data construction pipeline for generating high-quality ground-and-refer data for fine-grained document understanding. Additionally, we construct DOGR-Bench as the first comprehensive benchmark for evaluating the document grounding and referring capabilities of MLLMs. Furthermore, leveraging data generated by our engine, we develop DOGR, a pioneering MLLM that integrates text grounding and referring abilities into the dialogue and reasoning process. Our results show that DOGR can achieve versatile document grounding and referring while achieving promising performance on traditional document understanding tasks. We hope our work will facilitate practical document AI assistants.
8 Acknowledgement
This work was supported by the Early Career Scheme (No. CityU 21219323) and the General Research Fund (No. CityU 11220324) of the University Grants Committee (UGC), the NSFC Young Scientists Fund (No. 9240127), and the Donation for Research Projects (No. 9229164).
@@ B.2 Instruction Details
We add different response format prompts to different questions based on the format of the responses to help the model output corresponding results when users interact with it. For generated question and answer pairs, we combine different question and answer pairs with various response format prompts to obtain diverse grounded data. For Grounded Answering Data, we have 7 response format prompts to be added after the question, and the answer should be directly a simple text wrapped with “<ocr></ocr>” and followed by the coordinates wrapped with “<bbox></bbox>”. For Grounded Reasoning Data, we combine two random-chose prompts together to form the response format prompt, and the response should be such a sentence structure involves a segment of grounded reasoning followed by "Answer: " and a concise answer. For Grounded Open-ended Answering, we simply use the first part of the response format prompts for Grounded Reasoning, and the response should be a grounded reasoning sentence. For Plain text Answering, we add no response format prompt to keep consistent with the original question answering.
@@ Appendix C Qualitative Results
We present the inference results of DOGR for grounding, grounding-and-referring, and referring tasks. The specific coordinates of the annotations are omitted, and the grounded text is highlighted with colored boxes. The colors of the boxes within the document image correspond to the colors of the text boxes. For the grounding task, we present four examples. The first sample demonstrates that DOGR can perform fine-grained question answering on general document images with rich content and complex layouts, successfully grounding the corresponding information. The second sample illustrates the model’s excellent recognition and localization capabilities for diverse and small text in poster-type images. Additionally, we showcase two samples of grounding in chart-type images, which indicate that DOGR possesses a certain level of mathematical ability, enabling it to provide grounded reasoning during calculation, as well as the capability to estimate values in charts that lack textual annotations based on the axes. For the grounding-and-referring and referring tasks, DOGR is able to recognize the content of user-selected regions and provide reasonable grounded or plain textual reasoning and responses. DOGR exhibits robust fine-grained grounding and referring capabilities, allowing for reliable grounded reasoning and accommodating diverse user interactions, significantly enhancing the overall user experience.
@@ C.2 More Qualitative Results
C.2.1 DOGR-Bench Examples
As shown in Fig. 10, we present three inference examples of DOGR-Bench, demonstrating unlabeled value reading capability, grounding summarization capability, and referring summarization capability. The first figure is an example of reading unlabeled values. We add auxiliary mark boxes in red to the coordinate axes to facilitate reading the values, each box’s height represents 5. It can be observed that the model’s output values are accurate. The figure on the lower left is an example of grounding summarization for an entire document. DOGR is able to perceive the content of the entire document and perform grounding summarization effectively. The figure on the right is an example of referring summarization for a specific area. The model can also accomplish this task, thereby helping users improve reading efficiency.
Figure 10: DOGR’s inference results on DOGR-Bench.
@@ C.2.2 Other Examples
Strong generalization. As shown in the first row of examples, we feed DOGR with screenshots of the paper content and the specially shaped fan chart that the model does not process before. DOGR is able to correctly respond, demonstrating its strong generalization and usability in actual document reading scenarios.
Handwriting ground-and-refer ability. The middle sample showcases the model’s ability to recognize and grounding handwritten content. DOGR can fully understand this casual handwriting and provide grounded output. It is worth mentioning that our training data does not contain the grounding or referring tasks on such handwritten images.
Other capabilities. The bottom sample demonstrates an additional untrained capability of our model, specifically referring translation. DOGR is able to provide translations for the asked region. It is important to note that our training data do not include tasks similar to referring translation, even rarely includes languages other than English. Therefore, we believe that DOGR can effectively handle the relationship between regions and corresponding text, and seamlessly integrate the capabilities of large models with grounding and referring abilities.
@@ C.2.3 Failure Cases
Although DOGR demonstrates strong grounding and referring capabilities, there are still some shortcomings, as shown in Fig. 12.
Referring. The upper left figure illustrates an example of incorrect referring. It mistakenly associates the bounding box that should correspond to the text “INITIATIVE” in the question with the text “LOYALTY” below it. Additionally, its width is re-estimated according to the size corresponding to "LOYALTY," resulting in an incorrect answer.
Grounding. When encountering some unfamiliar text content such as tables in the upper right image, DOGR can understand the content, effectively identified the line breaks between “Doc” and “VQA” and merged them together, and provide correct answers, but the grounding boxes are inaccurate. There is also a issue of incomplete grounding content, such as the bottom sample, which often occurs when the content requiring grounding is interrupted or wrapped to the next line. Although this does not affect understanding, the text and bounding boxes provided by the model do not completely match.
@@
Figure 11: Other DOGR’s inference samples.
Figure 12: Failure cases of DOGR.
@@ D.1 Dataset Statistic
As shown in Tab. 6, we present the detailed composition of the DOGR dataset’s Multi-granular Parsing data and Instruction Tuning data. The Multi-granular Parsing data includes five granularities: word, phrase, line, paragraph, and full-page parsing. The Instruction-Tuning data comprises three types of grounded data: grounding referring, grounding-and-referring, and plain text Q&A. We provide the detailed data volume for each type.
@@
Table 6: The detailed composition of the DOGR dataset’s Multi-granular parsing data and Instruction-Tuning data.
@@ D.2 Construction Cost Details
In terms of multi-granular parsing data construction, DOGR-Engine can achieve boundary box annotation and content extraction on any document dataset with a similar rendering process. It can also scale up to a larger volume of PDF source files without additional manual costs. When constructing instruction fine-tuning data, our main cost lies in the API calls to GPT-4o. We use the version gpt-4o-2024-08-06. For poster and chart, our input is long text containing full page parsing data, while for PDF, the input is document images plus short prompt texts. During construction, we can generate multiple question and answer pairs for a single image simultaneously and use batch API requests to save costs. Ultimately, we can construct over 1,000 grounded question and answer pairs for an average cost of \$1, which is far more efficient and cost-effective than manual construction.
D.3 Hallucinations Avoidance
Regarding hallucinations, GPT-4o is quite capable of handling our tasks. By manual checking, we observe that hallucinations occur very rarely. We also filter out samples where the text and bounding boxes do not correspond correctly. Additionally, for DOGR-Bench, we manually filter out the samples that contain hallucinations or other errors.
@@ Appendix E Training Data Details
For the pre-training data, we utilize DocStruct4M [14] along with our 2.1M multi-granular parsing data to enhance DOGR’s foundational grounding and referring capabilities.
@@
DOGR- 83.34 60.96 41.67 36.95 55.73 49.66 58.62 50.34 47.27 51.47 68.86
DOGR 92.94 94.05 86.52 86.08 89.9 73.57 86.48 76.85 71.8 77.17 95.24
Appendix F Detailed Evaluation of MLLMs on DOGR-Bench.
F.1 Evaluation Prompt
Since models other than InternVL2.5 and QwenVL2.5 do not have a clearly defined format for grounding or referring content, we directly interpret the input using the prompt shown in Fig. 15(Right) and restrict the output format for evaluation. For InternVL2.5, the use of <ref></ref> is required to trigger grounding capabilities. Therefore, when outputs require answers and corresponding coordinates, we additionally include <ref></ref> to obtain a more standardized output for evaluation. For QwenVL2.5, since the model outputs pixel coordinates of the input image rather than the coordinates normalized according to the respective width and height, we first convert the coordinates in our input to facilitate the model’s accurate understanding of location information. We also restrict the output content to closely mimic the existing JSON format for evaluation purposes. For example, in reasoning tasks, the model first outputs a JSON list, then generates an explanation, and finally provides the answer. After obtaining the output, we normalize the output coordinates based on the input size for evaluation. The prompt we use for Qwen2.5 is shown in Fig. 15(Left).
F.2 Limited Grounding Ability
Although both InternVL2.5[7] and QwenVL2.5[1] can perform general image grounding, such as providing bounding boxes for a target based on a description, as shown in Fig. 16, when handling document grounding, we use a grounding-triggering prompt to obtain the output coordinates corresponding to the given content. InternVL2.5-72B consistently outputs inaccurate and unstable bounding boxes, while QwenVL2.5 also produces a large number of inaccurate bounding boxes. These simple examples illustrate the models’ shortcomings in document grounding capabilities.
F.3 IoU Threshold Sensitivity
We also conduct a statistical analysis of the $G_{a}$ $F1_{all}$ scores of different MLLMs at various IoU thresholds. As shown in Fig. 14, among the MLLMs analyzed, apart from DOGR, Qwen2.5VL-7B exhibits the best grounding capability. DOGR is not sensitive to changes in the IoU threshold, indicating that the limitation on DOGR’s $G_{a}$ $F1_{all}$ is due to the model’s reasoning capabilities. DOGR has already overcome the performance degradation caused by inaccurate grounding. In contrast, the insufficient basic grounding capabilities of other MLLMs make them very sensitive to changes in the IoU threshold, with their $G_{a}$ $F1_{all}$ scores dropping to nearly 0 when the IoU threshold is set to 0.5.
Figure 14: MLLM’s $G_{a}$ ${F1_{all}}$ comparison on different IoU threshold.
Figure 15: Prompts using in evaluation MLLMs on DOGR-Bench.
Figure 16: QwenVL2.5-72B and InternVL2.5-72B are all struggle with document grounding.
Appendix G More Experiments
@@
Effectiveness of our pretraining data. To better assess base performance of pretrained model using our multi-granular parsing data, we further introduce our text&bbox localization test set DOGR-Local15k comprising four granularities (word, phrase, line, paragraph) across three categories of data: poster, chart, and general document, similar to DocLocal4k[14]. DOGR- is pre-trained without using our constructed multi-granular parsing data. For the chart data, we claim that the chart type annotations in DocStruct4M for during DOGR-’s pre-training are inconsistent with our grounding training objectives, which aims to match the text and the bbox. In the original data, the text corresponds to bounding boxes of bar/line charts. Therefore, these value is not referenceable, and we mark them in gray. Moreover, in the case where the granularity of data categories for charts is relatively singular, we directly reported the results for “ALL”.
@@
As shown in Tab. 8, after incorporating our multi-granular parsing data for pre-training, the model exhibits enhancements in text recognition and grounding tasks. Our improvements on DOGR-Local15k are significant, thereby laying a solid foundation for accurate document grounding interactions.
The performance boost of data generated by GPT-4o. We utilize a weaker LLM, Gemini 1.5 flash, to generate 25k chart grounding QA data (Gemini25k). As shown in Tab.9, both Gemini25k and GPT-4o25k lead to improvement on ChartQA, which is mainly because constructed QA data enriches the original data. However, GPT-4o’s stronger instruction-following capability results in a larger amount of grounded texts in the constructed QA, making it better than gemini-1.5-flash on ${G_{r}}$ task.
Importance of accurate bounding box annotation. To validate importance of accurate bounding box annotation, we conduct experiments on chart data. We perform random offset and scaling of bboxes in GPT4o data within 30% to get 25k ${G_{r}}$ inaccurate training data GPT4o25k- with errors for chart. We compare the performance gain of 25k rerendered data with accurate bboxes and 25k grounded data with inaccurate bboxes on ChartQA and ${G_{r}}$ of chart on DO-Bench. Both text are from GPT-4o. As shown in Tab.10, adding GPT4o25k-, the performance on ChartQA is improved and the model possesses grounding capabilities. Adding accurate annotation data GPT4o25k- leads to more precise bboxes, thus largely improving the grounding performance(F1 of ${G_{r}}$).
Table 9: The performance boost of data generated by GPT-4o and weaker Gemini 1.5 flash.
Training Data ChartQA-Acc Gr-Acc Gr-F1
baseline(ChartQA) 80.92 - -
+GPT4o25k 81.88 67.0 32.91
+Gemini25k 81.98 50.12 1.10
Table 10: Accurate bbox is helpful for grounding.
Training Data ChartQA-Acc Gr-Acc Gr-F1
baseline(ChartQA) 80.92 - -
+GPT4o25k- 81.42 66.1 20.2
+GPT4o25k 81.88 67.0 32.91